# Source-Free Domain Adaptation with Diffusion-Guided Source Data Generation

### v1

@@ Abstract
This paper introduces a novel approach to leverage the generalizability capability of Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). Our proposed DM-SFDA method involves fine-tuning a pre-trained text-to-image diffusion model to generate source domain images using features from the target images to guide the diffusion process. Specifically, the pre-trained diffusion model is fine-tuned to generate source samples that minimize entropy and maximize confidence for the pre-trained source model. We then apply established unsupervised domain adaptation techniques to align the generated source images with target domain data. We validate our approach through comprehensive experiments across a range of datasets, including Office-31 (Saenko et al., 2010), Office-Home (Venkateswara et al., 2017), and VisDA (Peng et al., 2017). The results highlight significant improvements in SFDA performance, showcasing the potential of diffusion models in generating contextually relevant, domain-specific images.
@@ 1 Introduction
Deep Convolutional Neural Networks (CNNs) have demonstrated impressive performance on several visual tasks in recent years. However, the assumption that the distributions of the training and test sets are the same is crucial to the effectiveness of CNNs (Ding et al., 2022). Subsequently, a big drop in performance is generally observed when CNN-based models are deployed in real-world settings with a discrepancy in data distribution (Kundu et al., 2020). Domain Adaptation (DA) attempts to reduce this disparity to make these models perform well across multiple domains. Traditional DA approaches that rely on fixed source data might struggle to keep up with the pace of domain changes. Moreover, the rising prominence of data privacy regulations has led to a demand for DA techniques that can function effectively without relying on access to the source training data, a setting generally known as Source Free Domain Adaptation (SFDA).
@@
Most of the current state-of-the-art DA methods attain model adaptability by converging two disparate data distributions within a shared feature space, spanning both domains simultaneously (Ding et al., 2022). One way of achieving this in a source-free manner is to use synthetically generated source data. However, generating synthetic source data that accurately represents the diversity and complexity of the source domain can be difficult. Furthermore, if the synthetic data is not of high quality, it might introduce noise and inconsistencies, negatively impacting the model’s performance on the target domain. Notably, recent advancements in Diffusion Generative Models (DGMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated exceptional capabilities in producing diverse and high-quality images. Consequently, this paper aims to harness the generalizability of the state-of-the-art text-to-image diffusion models to the challenging task of SFDA.
@@
To address the challenges of data privacy and diversity in the reconstruction process, we present an innovative framework named Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). An overview of DM-SFDA is illustrated in Figure 1. The core idea of this approach is to use text-to-image diffusion models to generate images representative of the source domain based on the target domain and a pre-trained source network. Essentially, this involves fine-tuning a pre-trained text-to-image diffusion model to produce source samples that minimize the entropy for the pre-trained source model. The key contributions of our framework can be summarized as follows.
@@
The remainder of the paper is organized as follows. In Section 2, we discuss related work in the areas of DA, SFDA, and DGMs. In Section 3, we revisit the preliminary concepts that form the basis of our proposed approach. In Section 4, we discuss and formalize the problem definition and present our methodology. We present experimental results and analysis in Sections 5, 6, and 7. Finally, we conclude with a brief summary in Section 8.
@@ 2.1 Domain Adaptation
DA has its roots in (Ben-David et al., 2006), which focused on the role of good feature representations in their successful application for the task. The initial few works in DA adapted moment matching to align feature distributions between source and target domains (Long et al., 2015b; Huang et al., 2006; Long et al., 2016a; Sun et al., 2015; Tzeng et al., 2014). Subsequent works used adversarial learning-based approaches to tackle the problem of DA (Ganin & Lempitsky, 2014; Saito et al., 2017; Long et al., 2017; Ganin et al., 2016; Shu et al., 2018). Apart from this, many other techniques like (Xu et al., 2018; Tang et al., 2020a; Saito et al., 2020; Bermúdez-Chacón et al., 2020) have been proposed to tackle the task of DA.
@@ 2.2 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is a subtype of DA that aims to transfer knowledge from a labeled source domain to a different unlabeled target domain (Long et al., 2016b). Existing mainstream UDA methods can be categorized into two main types of methods: those that align the source and target domain distributions by designing specific metrics (Na et al., 2022, 2021; Li et al., 2021; Long et al., 2015a), and those that learn domain-invariant feature representations through adversarial learning (Westfechtel et al., 2024; Ganin & Lempitsky, 2015b; Long et al., 2018). However, the success of most of these methods depends on a huge amount of source data which might not be available in most practical scenarios.
@@ 2.3 Source-Free Domain Adaptation
SFDA has been considered in the literature as a means of reducing reliance on source data. As described in (Yu et al., 2023), the existing SFDA research can generally be categorized into two approaches: data-centric and model-centric. Model-centric methods employ techniques such as self-training and self-attention, while data-centric methods include domain-based reconstruction and image-based information extraction. Our proposed method follows the data-centric perspective to solve the SFDA task using source domain generation. 3C-GAN (Li et al., 2020) is a pioneering work in this area which uses a Generative Adversarial Network (GAN) to generate source-like images. Some other works like SDDA (Kurmi et al., 2021) and CPGA (Qiu et al., 2021) also solve the SFDA task using a similar approach.
@@ 2.4 DGMs for Domain Adaptation
Recently, there has been a significant shift in the landscape of generative modeling due to DGMs (Ho et al., 2020; Sohl-Dickstein et al., 2015), demonstrating impressive capabilities in generating highly realistic text-conditioned images. DGMs have also seen a growing interest in the DA community with many recent works using DGMs as input augmentation techniques. A recent example is a text-to-image diffusion model, employed by (Benigmim et al., 2023) to generate target domain images using source domain labels, thereby demonstrating the efficacy of diffusion models in One-Shot Unsupervised Domain Adaptation (OSUDA). DGMs, when trained on multiple source domains, have also been instrumental in guiding approximate inference in target domains, as reported by (Graikos et al., 2022). In our work, we leverage a recently introduced fine-tuning strategy for diffusion models called Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2023), to fine-tune the diffusion models using the output probability of the source model as an objective function.
@@ 3.1 Conditional Diffusion Probabilistic Models
3.2 Markov Decision Process and Reinforcement Learning
A Markov decision process (MDP) serves as a structured representation of problems involving a series of interconnected decisions. It is characterized by a set of components denoted as , where signifies the collection of possible states, stands for the available actions, signifies the initial state distribution, represents the transition pattern, and defines the reward mechanism. At each discrete time step denoted as , the agent observes a specific state represented as from the state space takes an action labeled as from the action space and in response receives a reward labeled as , subsequently transitioning to a novel state drawn from the distribution . The agent’s decision-making is guided by a policy that dictates actions based on states. During the agent’s engagement with the MDP, it generates trajectories, which are sequences comprising both states and actions, conventionally presented as . Within the context of reinforcement learning (RL), the primary aim of the agent is to maximize a metric termed , representing the expected cumulative reward garnered over trajectories sampled in accordance with its policy:
3.3 Denoising as a multi-step MDP
When the sampler is fixed, the denoising process can be reframed as a multi-step MDP (Black et al., 2023). The iterative denoising process can be mapped to the following MDP:
Here, is the Dirac delta distribution with a non-zero density only at .
@@ 4 Proposed Method
4.1 Notations
Following the notations used in (Yu et al., 2023), in this paper, we represent a domain as . Each domain consists of a dataset and an associated label set . A dataset comprises of an instance set , derived from a -dimensional marginal distribution , and a label set where represents the total number of samples and represents the total number of classes.
4.2 Problem Definition
@@ 4.3 Method Overview
An overview of the proposed method DM-SFDA for source free domain adaptation using text-to-image diffusion models can be seen in Figure 1. The key idea behind the approach is to leverage the generalizability of the state-of-the-art text-to-image diffusion models to tackle the task of SFDA. As shown in Figure 2, the training pipeline contains the following four phases: I) Selective Pseudo Labeling Target Data II) Fine-tuning Diffusion Model on Target Data, III) Source Data Generation using DDPO, IV) Unsupervised Domain Adaptation. The first phase involves selective pseudo labeling target data using the pre-trained source model. The second phase involves jointly fine-tuning an implicit image captioner along with a pre-trained text-to-image diffusion model on the target images. Subsequently, the pre-trained source model is used to fine-tune this diffusion model using Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2023) to generate Source Images. Finally, the generated source images and the target images are used to perform unsupervised domain adaptation. Each of these phases are described in detail in the following subsections.
@@ 4.4 Phase I: Selective Pseudo Labeling Target Data
The proposed DM-SFDA pipeline requires labeled target data to generate data from the source domain. Therefore, the initial phase of our proposed pipeline addresses the challenge of unlabeled target data by selecting reliable labels for the target samples using a selective pseudo-labeling strategy. This approach is akin to the one proposed in (Karim et al., 2023). As shown in (Guillory et al., 2021), prediction confidence and difference in entropy can be reliable measures of pseudo-label accuracy and estimate different types of domain shifts. Therefore, prediction confidence and average entropy are used as metrics to assess label reliability. A binary reliability score () is assigned to each sample in the target data, determined by their prediction confidence and prediction uncertainty (), as illustrated below
Here, and represent the selection thresholds for confidence and uncertainity. Taking the average as a threshold eliminates the requirement of per-dataset hyper-parameter tuning and makes the selection process highly adaptive. Furthermore, aleatoric uncertainty (Hüllermeier & Waegeman, 2021) is used since it better addresses the concern of domain shift.
@@ 4.5 Phase II: Diffusion Model Finetuning on Target Data
The second phase of the proposed DM-SFDA pipeline involves fine-tuning a text-to-image diffusion model on the target data. However, the unavailability of labels classes poses a challenge as we don’t have any textual cues to guide the diffusion process. Furthermore, passing an empty caption would also be sub-optimal (Xu et al., 2023). This issue of unavailability of target labels is tackled by using an implicit image captioner, as used in (Xu et al., 2023). Instead of using an off-the-shelf captioning network to generate captions, a network is trained to generate an implicit text embedding from the input image itself. The orange block in Figure 2 shows the architecture of the implicit image captioner. Specifically, to derive the implicit text embedding of the image, a pre-trained frozen image encoder, (e.g. from CLIP (Radford et al., 2021) is used to encode the input image into its embedding space. A learned MLP is added to project the image embedding into an implicit text embedding. The image captioner is jointly finetuned along with the diffusion model to generate implicit text embedding for the input target image. During the training phase, the parameters of the image encoder is unchanged and we only finetune the parameters of the MLP and UNet. Formally, the textual features of a target image are extracted as follows:
4.6 Phase III: Source Data Generation using DDPO
In the third phase, we use the method proposed in (Black et al., 2023) to further fine-tune our diffusion model to generate source-like images. As explained in (Black et al., 2023), when we formulate the denoising process of a diffusion model as a multi-step MDP, and we use a standard sampler parameterized as in Equation 2, the policy becomes an isotropic Gaussian. This simplification allows for the evaluation of exact action likelihoods and gradients of these likelihoods with respect to the diffusion model parameters. Subsequently, with access to likelihood and likelihood gradients, the fine-tuning alternates collecting trajectories via sampling and updating parameters via gradient ascent on the diffusion RL objective :
The pre-trained source model serves as the reward function for the fine-tuning procedure. The ultimate objective is to maximize the average classification confidence and minimize the average entropy on the generated source data. The reward function used for fine-tuning the model is shown below:
After the completion of the DDPO fine-tuning, the diffusion model produces source images by utilizing the implicit target image captions as guiding prompts.
@@ 4.7 Phase IV: Unsupervised Domain Adaptation
The fourth phase starts by labeling the generated source domain images using the pre-trained source model. Once we have the reconstructed and labeled source domain data, we have effectively converted the initial problem of SFDA to a standard Unsupervised Domain Adaptation (UDA) problem which can be solved using any of the pre-existing approaches. Since we chose a ResNet backbone as the pre-trained source model, we experiment with the current state-of-the-art ResNet-based UDA approaches. For the purpose of this paper, we experiment with the following approaches:
4.7.1 Bridging Domain Spaces (FixBi)
FixBi (Na et al., 2021) tackles the problem of domain adaptation by compensating for the large discrepancy between the source and target domains. This is done by generating multiple intermediate augmented domains by using a fix ration-based mixup data augmentation approach. Complementary models are trained on each of these intermediate domains that teach each other to bridge the domain gap by using a confidence-based learning where one model teaches the other model using the positive pseudo-labels or teach itself using the negative pseudo-labels. Through the confidence-based learning approach, the two models with different characteristics gradually get closer to the target domain. Furthermore, a consistency regularization is used to ensure a stable convergence of training both models.
@@ 4.7.1 Bridging Domain Spaces (FixBi)
| ResNet-50 (He et al., 2015) | ✗ | 68.9 | 68.4 | 62.5 | 96.7 | 60.7 | 99.3 | 76.1 | ||
| MCC (Jin et al., 2019) | ✗ | 95.6 | 95.4 | 72.6 | 98.6 | 73.9 | 100.0 | 89.4 | ||
| GSDA (Hu et al., 2020) | ✗ | 94.8 | 95.7 | 73.5 | 99.1 | 74.9 | 100.0 | 89.7 | ||
| SRDC (Tang et al., 2020b) | ✗ | 95.8 | 95.7 | 76.7 | 99.2 | 77.1 | 100.0 | 90.8 | ||
| FixBi (Na et al., 2021) | ✗ | 95.0 | 96.1 | 78.7 | 99.3 | 79.4 | 100.0 | 91.4 | ||
| CoVi (Na et al., 2022) | ✗ | 98.0 | 97.6 | 77.5 | 99.3 | 78.4 | 100.0 | 91.8 | ||
| SHOT (Liang et al., 2020) | ✓ | 94.0 | 90.1 | 74.7 | 98.4 | 74.3 | 99.9 | 88.6 | ||
| 3C-GAN (Li et al., 2020) | ✓ | 92.7 | 93.7 | 75.3 | 98.5 | 77.8 | 99.8 | 89.6 | ||
| NRC (Yang et al., 2021) | ✓ | 96.0 | 90.8 | 75.3 | 99.0 | 75.0 | 100.0 | 89.4 | ||
| NRC++ (Yang et al., 2023) | ✓ | 95.9 | 91.2 | 75.5 | 99.1 | 75.0 | 100.0 | 89.5 | ||
| DM-SFDA (FixBi) | ✓ | 93.0 | 93.5 | 77.4 | 98.7 | 78.3 | 99.7 | 90.1 | ||
| DM-SFDA (CoVi) | ✓ | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
@@
| ResNet-50 (He et al., 2015) | ✗ | 34.9 | 50.0 | 58.0 | 37.4 | 41.9 | 46.2 | 38.5 | 31.2 | 60.4 | 53.9 | 41.2 | 59.9 | 46.1 |
| MCC (Jin et al., 2019) | ✗ | 88.1 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA (Hu et al., 2020) | ✗ | 61.3 | 76.1 | 79.4 | 65.4 | 73.3 | 74.3 | 65.0 | 53.2 | 80.0 | 72.2 | 60.6 | 83.1 | 70.3 |
| SRDC (Tang et al., 2020b) | ✗ | 52.3 | 76.3 | 81.0 | 69.5 | 76.2 | 78.0 | 68.7 | 53.8 | 81.7 | 76.3 | 57.1 | 85.0 | 71.3 |
| FixBi (Na et al., 2021) | ✗ | 58.1 | 77.3 | 80.4 | 67.7 | 79.5 | 78.1 | 65.8 | 57.9 | 81.7 | 76.4 | 62.9 | 86.7 | 72.7 |
| CoVi (Na et al., 2022) | ✗ | 58.5 | 78.1 | 80.0 | 68.1 | 80.0 | 77.0 | 66.4 | 60.2 | 82.1 | 76.6 | 63.6 | 86.5 | 73.1 |
| SHOT (Liang et al., 2020) | ✓ | 57.1 | 78.1 | 81.5 | 68.0 | 78.2 | 78.1 | 67.4 | 54.9 | 82.2 | 73.3 | 58.8 | 84.3 | 71.8 |
| NRC (Yang et al., 2021) | ✓ | 57.7 | 80.3 | 82.0 | 68.1 | 79.8 | 78.6 | 65.3 | 56.4 | 83.0 | 71.0 | 58.6 | 85.6 | 72.2 |
| NRC++ (Yang et al., 2023) | ✓ | 57.8 | 80.4 | 81.6 | 69.0 | 80.3 | 79.5 | 65.6 | 57.0 | 83.2 | 72.3 | 59.6 | 85.7 | 72.5 |
| DM-SFDA (FixBi) | ✓ | 57.9 | 77.1 | 80.1 | 67.3 | 79.4 | 78.0 | 65.7 | 57.5 | 81.6 | 75.3 | 62.7 | 86.5 | 72.4 |
| DM-SFDA (CoVi) | ✓ | 58.0 | 77.8 | 79.6 | 67.9 | 80.0 | 76.8 | 66.0 | 58.8 | 82.0 | 75.8 | 63.0 | 86.3 | 72.6 |

### v2

@@ Abstract
This paper introduces a novel approach to leverage the generalizability of Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). Our proposed DM-SFDA method involves fine-tuning a pre-trained text-to-image diffusion model to generate source domain images using features from the target images to guide the diffusion process. Specifically, the pre-trained diffusion model is fine-tuned to generate source samples that minimize entropy and maximize confidence for the pre-trained source model. We then use a diffusion model-based image mixup strategy to bridge the domain gap between the source and target domains. We validate our approach through comprehensive experiments across a range of datasets, including Office-31 [39], Office-Home [48], and VisDA [35]. The results demonstrate significant improvements in SFDA performance, highlighting the potential of diffusion models in generating contextually relevant, domain-specific images.
@@ 1 Introduction
Deep Convolutional Neural Networks (CNNs) have demonstrated impressive performance on several visual tasks in recent years. However, the assumption that the distributions of the training and test sets are the same is crucial to the effectiveness of CNNs [5]. Subsequently, a big drop in performance is generally observed when CNN-based models are deployed in real-world settings with a discrepancy in data distribution [21]. Domain Adaptation (DA) attempts to reduce this disparity to make these models perform well across multiple domains. Traditional DA approaches that rely on fixed source data might struggle to keep up with the pace of domain changes. Moreover, the rising prominence of data privacy regulations has led to a demand for DA techniques that can function effectively without relying on access to the source training data, a setting generally known as Source Free Domain Adaptation (SFDA).
@@
Most of the current state-of-the-art DA methods attain model adaptability by converging two disparate data distributions within a shared feature space, spanning both domains simultaneously [5]. One way of achieving this in a source-free manner is to use synthetically generated source data. However, generating synthetic source data that accurately represents the diversity and complexity of the source domain can be difficult. Furthermore, if the synthetic data is not of high quality, it might introduce noise and inconsistencies, negatively impacting the model’s performance on the target domain. Notably, recent advancements in Diffusion Generative Models (DGMs) [15, 43] have demonstrated exceptional capabilities in producing diverse and high-quality images. Consequently, this paper aims to harness the generalizability of the state-of-the-art text-to-image diffusion models to the challenging task of SFDA.
@@
To address the challenges of data privacy and diversity in the reconstruction process, we present an innovative framework named Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). An overview of DM-SFDA is illustrated in Figure 1. The core idea of this approach is to use text-to-image diffusion models to generate images representative of the source domain based on the target domain and a pre-trained source network. Essentially, this involves fine-tuning a pre-trained text-to-image diffusion model to produce source samples that minimize the entropy for the pre-trained source model. The key contributions of our framework can be summarized as follows:
1.
We propose a novel framework to enhance model performance in unseen domains while simultaneously addressing the challenges posed by limited access to source data and the increasing emphasis on data privacy.
2.
Our novel framework harnesses the generalization capabilities of Diffusion Models to improve the data diversity and completeness within the reconstructed source data.
3.
Through extensive qualitative and quantitative analyses of several traditional and state-of-the-art baselines and in-depth analysis, we demonstrate the effectiveness of the proposed pipeline.
@@
The remainder of the paper is organized as follows. In Section 2, we discuss related work in the areas of DA, SFDA, and DGMs. In Section 3, we revisit the preliminary concepts that form the basis of our proposed approach. In Section 4, we discuss and formalize the problem definition and present our methodology. We present experimental results and analysis in Sections 5 and 6. Finally, we discuss the limitations of our work in Section 7 and conclude with a brief summary in Section 8.
@@ 2.1 Domain Adaptation
DA has its roots in [1], which focused on the role of good feature representations in their successful application for the task. The initial few works in DA adapted moment matching to align feature distributions between source and target domains [28, 17, 29, 44, 47]. Subsequent works used adversarial learning-based approaches to tackle the problem of DA [7, 40, 31, 10, 42]. Apart from this, many other techniques like [50, 45, 41, 3] have been proposed to tackle the task of DA.
@@ 2.2 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is a subtype of DA that aims to transfer knowledge from a labeled source domain to a different unlabeled target domain [30]. Existing mainstream UDA methods can be categorized into two main types of methods: those that align the source and target domain distributions by designing specific metrics [34, 33, 24, 27], and those that learn domain-invariant feature representations through adversarial learning [49, 9, 32]. However, the success of most of these methods depends on a huge amount of source data which might not be available in most practical scenarios.
@@ 2.3 Source-Free Domain Adaptation
SFDA has been considered in the literature as a means of reducing reliance on source data. As described in [55], the existing SFDA research can generally be categorized into two approaches: data-centric and model-centric. Model-centric methods employ techniques such as self-training and self-attention, while data-centric methods include domain-based reconstruction and image-based information extraction. Our proposed method follows the data-centric perspective to solve the SFDA task using source domain generation. 3C-GAN [25] is a pioneering work in this area which uses a Generative Adversarial Network (GAN) to generate target-like images and simulatneously adapts the source pre-trained model. Some other works like SDDA [22] and CPGA [37] also solve the SFDA task using a similar data-generation approach. More recently, other works like AaD [52] and Co-learn [57] have also shown state-of-the-art performance across various datasets and tasks.
@@ 2.4 DGMs for Domain Adaptation
Recently, there has been a significant shift in the landscape of generative modeling due to DGMs [15, 43], demonstrating impressive capabilities in generating highly realistic text-conditioned images. DGMs have also seen a growing interest in the DA community with many recent works using DGMs as input augmentation techniques. A recent example is a text-to-image diffusion model, employed by [2] to generate target domain images using source domain labels, thereby demonstrating the efficacy of diffusion models in One-Shot Unsupervised Domain Adaptation (OSUDA). DGMs, when trained on multiple source domains, have also been instrumental in guiding approximate inference in target domains, as reported by [11]. In our work, we leverage a recently introduced fine-tuning strategy for diffusion models called AlignProp [36], to fine-tune the diffusion models using the output probability of the source model as an objective function.
@@ 3.1 Conditional Diffusion Probabilistic Models
3.2 Markov Decision Process
A Markov decision process (MDP) serves as a structured representation of problems involving a series of interconnected decisions. It is characterized by a set of components denoted as , where signifies the collection of possible states, stands for the available actions, signifies the initial state distribution, represents the transition pattern, and defines the reward mechanism. At each discrete time step denoted as , the agent observes a specific state represented as from the state space takes an action labeled as from the action space and in response receives a reward labeled as , subsequently transitioning to a novel state drawn from the distribution . The agent’s decision-making is guided by a policy that dictates actions based on states. During the agent’s engagement with the MDP, it generates trajectories, which are sequences comprising both states and actions, conventionally presented as .
@@ 4 Proposed Method
4.1 Notations and Problem Definition
Following the notations used in [55], in this paper, we represent a domain as . Each domain consists of a dataset and an associated label set . A dataset comprises of an instance set , derived from a -dimensional marginal distribution , and a label set where represents the total number of samples and represents the total number of classes.
@@ 4.2 Method Overview
An overview of the proposed method DM-SFDA for source free domain adaptation using text-to-image diffusion models can be seen in Figure 1. The key idea behind the approach is to leverage the generalizability of the state-of-the-art text-to-image diffusion models to tackle the task of SFDA. The training pipeline contains the following four phases: I) Selective Pseudo Labeling Target Data II) Fine-tuning Diffusion Model on Target Data, III) Source Data Generation using AlignProp, IV) Unsupervised Domain Adaptation. The first phase involves selective pseudo labeling target data using the pre-trained source model. The second phase involves finetuning a pre-trained diffusion model to learn concepts in the target domain using Textual Inversion [6]. Subsequently, the pre-trained source model is used to fine-tune this diffusion model using AlignProp [36] to generate Source Images. Finally, a diffusion model-based domain mixup strategy is used to perform unsupervised domain adaptation. Each of these phases are described in detail in the following subsections.
@@ 4.3 Phase I: Selective Pseudo Labeling Target Data
The proposed DM-SFDA pipeline requires labeled target data to generate data from the source domain. Therefore, the initial phase of our proposed pipeline addresses the challenge of unlabeled target data by selecting reliable labels for the target samples using a selective pseudo-labeling strategy. This approach is akin to the one proposed in [20]. As shown in [13], prediction confidence and difference in entropy can be reliable measures of pseudo-label accuracy and estimate different types of domain shifts. Therefore, prediction confidence and average entropy are used as metrics to assess label reliability. A binary reliability score () is assigned to each sample in the target data, determined by their prediction confidence and prediction uncertainty (), as illustrated below:
Here, and represent the selection thresholds for confidence and uncertainity. Taking the average as a threshold eliminates the requirement of per-dataset hyper-parameter tuning and makes the selection process highly adaptive. Furthermore, aleatoric uncertainty [18] is used since it better addresses the concern of domain shift.
@@ 4.4 Phase II: Diffusion Model Finetuning on Target Data
The second phase of the proposed DM-SFDA pipeline involves fine-tuning a text-to-image diffusion model on the target data using LoRA. However, the lack of class labels poses a challenge as there are no textual cues to guide the diffusion process. To address this, we employ a recently introduced fine-tuning strategy called Textual Inversion [6]. Using images from the target domain, Textual Inversion learns to represent objects in the images through new "words" in the embedding space of a pre-trained text-to-image model. As illustrated in Figure 1, we assign a placeholder string "<class-{idx}>" for the newly learned concepts, using class indices from the selective pseudo labeling done in Phase I. These new identifiers are then used as textual cues to guide the diffusion process in subsequent phases of the pipeline.
4.5 Phase III: Source Data Generation using AlignProp
In the third phase, we use the method proposed in [36] to further fine-tune our diffusion model to generate source-like images. This finetuning is done by transforming the denoising process of a diffusion model into a differentiable recurrent policy. The iterative denoising process is mapped to the following single-step MDP:
Here, is the reward function which is dependent on the generated images. In our case, we define the reward function to use the information in the pre-trained source model to guide the diffusion process to generate source-like images. Inspired by the loss function used in DAFL [4], our reward function consists of the following three components to extract maximum information from the source model:
Confidence Reward: The confidence reward function makes sure that higher confidence predictions are assigned a higher reward.
(5) -
One-Hot Reward: As described in [4], the outputs of the source-model should be similar to the training data if the input follows the training distribution. Therefore, the one-hot reward function assigns higer rewards to the samples that generate one-hot like predictions.
(6) -
Batch Norm Statistics (BNS) Reward: As described in [54], to effectively match the low level and high level feature maps of the generated images, we can match the Batch Norm Statistics of the pre-trained neural network with the generated images. The BNS reward function ensures that a higer reward is given to the samples which have similar BNS as that of the model.
The final reward function used for finetuning the diffusion model is defined as:
Using the above state, action and reward functions, and the class prompts , the parameters of the diffusion model are updated using gradient descent on the following loss function:
The training is done using backpropagation through time on the recurrent policy . Furthermore, as described in [36], in order to reduce the memory overload, the LoRA weights are finetuned along with using a truncated backpropagation through time (TBTT) instead of a full backpropagation through time (BPTT). After the completion of the AlignProp fine-tuning, the diffusion model produces source images by utilizing the "<class-{idx}>" placeholder as guiding prompt.
@@ 4.6 Phase IV: Unsupervised Domain Adaptation
The fourth phase starts by labeling the generated source domain images using the pre-trained source model. Once we have the reconstructed and labeled source domain data, we have effectively converted the initial problem of SFDA to a standard Unsupervised Domain Adaptation (UDA) problem. Since, all of the diffusion model finetuning happens via Low-Rank Adapters, we can use the finetuned weight patches to effectively compensate the domain discrepancy between the source and target domain. This is done by generating multiple intermediate augmented domains by altering the scaling parameter while applying the LoRA patch to the pretrained model. Figure 2, shows the visualization of the diffusion model-based inter-domain mixup. Subsequently, we use the approach proposed in [33] to train two complementary models on each of these intermediate domains that teach each other to bridge the domain gap by using a confidence-based learning where one model teaches the other model using the positive pseudo-labels or teach itself using the negative pseudo-labels. Through the confidence-based learning approach, the two models with different characteristics gradually get closer to the target domain. Furthermore, a consistency regularization is used to ensure a stable convergence of training both models.
@@
| ResNet-50 [14] | ✗ | 68.9 | 68.4 | 62.5 | 96.7 | 60.7 | 99.3 | 76.1 | ||
| MCC [19] | ✗ | 95.6 | 95.4 | 72.6 | 98.6 | 73.9 | 100.0 | 89.4 | ||
| GSDA [16] | ✗ | 94.8 | 95.7 | 73.5 | 99.1 | 74.9 | 100.0 | 89.7 | ||
| SRDC [46] | ✗ | 95.8 | 95.7 | 76.7 | 99.2 | 77.1 | 100.0 | 90.8 | ||
| FixBi [33] | ✗ | 95.0 | 96.1 | 78.7 | 99.3 | 79.4 | 100.0 | 91.4 | ||
| CoVi [34] | ✗ | 98.0 | 97.6 | 77.5 | 99.3 | 78.4 | 100.0 | 91.8 | ||
| ICON [56] | ✗ | 97.0 | 93.3 | 79.4 | 99.2 | 78.3 | 100.0 | 91.2 | ||
| SHOT [26] | ✓ | 94.0 | 90.1 | 74.7 | 98.4 | 74.3 | 99.9 | 88.6 | ||
| 3C-GAN [25] | ✓ | 92.7 | 93.7 | 75.3 | 98.5 | 77.8 | 99.8 | 89.6 | ||
| NRC [51] | ✓ | 96.0 | 90.8 | 75.3 | 99.0 | 75.0 | 100.0 | 89.4 | ||
| NRC++ [53] | ✓ | 95.9 | 91.2 | 75.5 | 99.1 | 75.0 | 100.0 | 89.5 | ||
| AaD [52] | ✓ | 96.4 | 92.1 | 75.0 | 99.1 | 76.5 | 100.0 | 89.9 | ||
| AaD w/ Co-learn [57] | ✓ | 97.6 | 98.7 | 82.1 | 99.3 | 80.1 | 100.0 | 93.0 | ||
| DM-SFDA | ✓ | 97.7 | 99.0 | 82.7 | 99.3 | 83.5 | 100.0 | 93.7 |
@@
| ResNet-50 [14] | ✗ | 34.9 | 50.0 | 58.0 | 37.4 | 41.9 | 46.2 | 38.5 | 31.2 | 60.4 | 53.9 | 41.2 | 59.9 | 46.1 |
| MCC [19] | ✗ | 88.1 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA [16] | ✗ | 61.3 | 76.1 | 79.4 | 65.4 | 73.3 | 74.3 | 65.0 | 53.2 | 80.0 | 72.2 | 60.6 | 83.1 | 70.3 |
| SRDC [46] | ✗ | 52.3 | 76.3 | 81.0 | 69.5 | 76.2 | 78.0 | 68.7 | 53.8 | 81.7 | 76.3 | 57.1 | 85.0 | 71.3 |
| FixBi [33] | ✗ | 58.1 | 77.3 | 80.4 | 67.7 | 79.5 | 78.1 | 65.8 | 57.9 | 81.7 | 76.4 | 62.9 | 86.7 | 72.7 |
| CoVi [34] | ✗ | 58.5 | 78.1 | 80.0 | 68.1 | 80.0 | 77.0 | 66.4 | 60.2 | 82.1 | 76.6 | 63.6 | 86.5 | 73.1 |
| ICON [56] | ✗ | 63.3 | 81.3 | 84.5 | 70.3 | 82.1 | 81.0 | 70.3 | 61.8 | 83.7 | 75.6 | 68.6 | 87.3 | 75.8 |
| SHOT [26] | ✓ | 57.1 | 78.1 | 81.5 | 68.0 | 78.2 | 78.1 | 67.4 | 54.9 | 82.2 | 73.3 | 58.8 | 84.3 | 71.8 |
| NRC [51] | ✓ | 57.7 | 80.3 | 82.0 | 68.1 | 79.8 | 78.6 | 65.3 | 56.4 | 83.0 | 71.0 | 58.6 | 85.6 | 72.2 |
| NRC++ [53] | ✓ | 57.8 | 80.4 | 81.6 | 69.0 | 80.3 | 79.5 | 65.6 | 57.0 | 83.2 | 72.3 | 59.6 | 85.7 | 72.5 |
| AaD [52] | ✓ | 59.3 | 79.3 | 82.1 | 68.9 | 79.8 | 79.5 | 67.2 | 57.4 | 83.1 | 72.1 | 58.5 | 85.4 | 72.7 |
| AaD w/ Co-learn [57] | ✓ | 65.1 | 86.0 | 87.0 | 76.8 | 86.3 | 86.5 | 74.4 | 66.1 | 87.7 | 77.9 | 66.1 | 88.4 | 79.0 |
| DM-SFDA | ✓ | 68.5 | 89.6 | 83.3 | 70.0 | 85.8 | 87.4 | 71.3 | 69.6 | 88.2 | 77.8 | 68.5 | 88.7 | 79.5 |